A researcher had conducted an ANOVA and found that at least one of the following methods of studying for an exam:

1) office hours regularly,

2) both office hours and review sessions,

3) just review sessions, and

4) neither office hours nor review sessions.

differed significantly from the other methods in terms of stats exam scores [F(3, 16) = 9.84, p< .05]. Below is the original ANOVA source table:

Source             SS                    DF                   MS                   F

Between              85                3                      28.33 9.84    

Within             46                    16                    2.88

The researcher wanted to conduct pairwise comparisons to find which of the methods differed from the others in resulting stats exam scores. The researcher’s raw data were as follows:

Office

(O)

Office & Review (O&R)

Review

(R)

None

(N)

12

12

8

6

11

10

7

9

14

15

10

11

11

14

9

10

12

14

6

9

X̄=12 SS=6

X̄=13 SS=16

X̄=8 SS=10

X̄=9 SS=14

Conduct pairwise comparisons via both Tukey’s HSD and Scheffe’s test. Show all work
Answer #1

SOLUTION

(1) Ho: All four groups have the same mean (μ1 = μ2 = μ3 = μ4)

Ha: At least two of the four groups have significantly different means

(2) α = 0.05, dfB = 4 - 1 = 3, dfT = 20 - 1 = 19, dfW = 19 - 3 = 16

(3 and 4)

Group 1 Group 2 Group 3 Group 4 12 12 15 14 14 10 12 Sum- 65 45 Average- 12 726 1.225 419 1.871 14 330 1.581 10 861 St. Dev.-

The total sample size is N 20. Therefore, the total degrees of freedom are: dfrotal 20 - 1-19 Also, the between-groups degree

The within sum of squares is computed as shown in the calculation below SSuithin S Suithingroup 6+16+10+14 46 The between sum

(1) Null and Alternative Hypotheses The following null and altermative hypotheses need to be tested Ha: Not all means are equ

(5 Conclusion It is concluded that the null hypothesis Ho is rejected. Therefore, there is enough evidence to claim that not